### v1

# Balancing Act: Distribution-Guided Debiasing in Diffusion Models

Abstract
Diffusion Models (DMs) have emerged as powerful generative models with unprecedented image generation capability. These models are widely used for data augmentation and creative applications. However, DMs reflect the biases present in the training datasets. This is especially concerning in the context of faces, where the DM prefers one demographic subgroup vs others (eg. female vs male). In this work, we present a method for debiasing DMs without relying on additional data or model retraining. Specifically, we propose Distribution Guidance, which enforces the generated images to follow the prescribed attribute distribution. To realize this, we build on the key insight that the latent features of denoising UNet hold rich demographic semantics, and the same can be leveraged to guide debiased generation. We train Attribute Distribution Predictor (ADP) - a small mlp that maps the latent features to the distribution of attributes. ADP is trained with pseudo labels generated from existing attribute classifiers. The proposed Distribution Guidance with ADP enables us to do fair generation. Our method reduces bias across single/multiple attributes and outperforms the baseline by a significant margin for unconditional and text-conditional diffusion models. Further, we present a downstream task of training a fair attribute classifier by rebalancing the training set with our generated data. Project Page
1 Introduction
Figure 1: a) Random sampling from a pretrained DM [33] generates images with gender imbalance. b) Proposed method takes a user-defined reference attribute $\mathbf{p^{a}_{ref}}$ and performs distribution guidance on a pretrained DM. c) Sampling with distribution guidance results in fair generation that follow user define $\mathbf{p^{a}_{ref}}$.
Recent advancements in Diffusion Models (DM) [11, 33, 42] have garnered much interest among researchers in evaluating the quality of the generated content. These models are not only used to generate realistic content but also to augment real datasets [45, 24] for downstream tasks. However, these popular models have been found to synthesize biased content with respect to multiple demographic factors like gender, race, age, etc., which can be detrimental to society once these models are deployed in the real world [29, 21, 34]. The problem is largely caused by the images used to train these models, as the outputs of these models are governed by these training datasets [29, 22]. Effects of such harmful biases have been shown by multiple recent works involving studies on not only DMs [29, 21, 34], but also GANs and other generative models [22, 5, 12]. In fact, Perera et al. [29] show that unconditional DMs–even when trained with balanced data–amplify racial biases, leading to the generation of more white-skinned faces than dark-skinned ones. Biases in the generated data are even more evident in large text-to-image DMs, e.g., models mostly tend to generate a specific gender with a given profession (like male and doctor) [23, 52, 39].
Existing works on debiasing, either require a reference dataset  [5, 52] and/or allow retraining of the model [5, 50, 51]. On the contrary, our work aims to mitigate biases in both unconditional and conditional DMs, enabling fair generation without model retraining. We propose a setting where we provide a reference attribute distribution ($\mathbf{p^{a}_{ref}}$) for sensitive attribute $a$ and query the DM to sample images following $\mathbf{p^{a}_{ref}}$. This framework allows for adapting any existing DM to a pre-defined distribution. E.g., a user can provide $\mathbf{p^{a}_{ref}}$ as a uniform histogram of sensitive attributes to generate balanced attribute distribution. We believe defining $\mathbf{p^{a}_{ref}}$ provides just enough information to condition the DM for fair generation. This is an extremely practical setting for debiasing and is particularly important for large text-to-image DMs [33, 36, 32] where retraining or fine-tuning is computationally intensive.
One plausible approach for fair generation is to guide every generated sample with attribute classifiers following classifier guidance [8]. However, such a framework, though simple, is overly restrictive as it requires presetting and enforcing attributes for each sample individually (which we call sample guidance). Such constraints during denoising result in inferior generation quality, as discussed in Sec.4.1. Instead, we propose to jointly denoise a whole batch of samples and guide the process with $\mathbf{p^{a}_{ref}}$ (which we call distribution guidance). Specifically, we push the generated batch attribute distribution $\mathbf{{p^{a}_{\theta}}}$ and $\mathbf{p^{a}_{ref}}$ close to each other during the reverse process. Distribution guidance provides more flexibility to each sample during generation as it does not enforce a preset of attributes on a sample basis. This results in fair generation without sacrificing the generation quality.
A major challenge for guidance-based conditioning is that it requires separate image classifiers for each noise scale of the diffusion process. To overcome this, we propose to perform guidance in semantically rich features space - h-space [18] of DMs. Specifically, we train an Attribute Distribution Predictor (ADP) that predicts attribute distribution directly from the h-space features. As ADP is trained on rich and discriminative h-space features, it is - a) implemented as a linear layer, b) requires minimal training data, and c) fast training and inference. Finally, during inference, we steer the h-space representation by matching the predictions from ADP to $\mathbf{p^{a}_{ref}}$.
We extensively evaluate our proposed method for the fair generation of single and multi-attribute cases for face generation. Additionally, we present the results of our method on Stable Diffusion [33], a large text-to-image DM. Further, as downstream application train debiased attribute classifiers by augmenting the training data for minority subgroups.
The major contributions of this work are the following:
  1. 1.
    A novel setting for debiasing existing DMs without retraining given a reference attribute distribution.
  2. 2.
    Distribution guidance to condition the reverse diffusion process on a reference attribute distribution.
  3. 3.
    Propose guidance in the intermediate features of diffusion network (h-space), which leads to data-efficient training and fast generation.
2 Related Works
Biases in Generative Models. While generative models like Generative Adversarial Networks and Diffusion Models have become the de-facto tools for image generation in recent times, studies show that they are not free of biases [29, 21, 34, 22, 13]. Perera et al.  [29] show that unconditional diffusion models amplify the biases in the training data with respect to gender, race, and age. Luccioni et al.  [21] identify and quantify social biases in images generated by popular text-to-image models like DALL-E 2, and Stable Diffusion v1.4 and 2. Maluleke et al.  [22] study racial biases in GANs and find that GANs mimic the racial distribution in the training data.
Debiasing generative models by retraining. This line of work focuses on mitigating biases in generative models by retraining them [27, 50, 47, 38, 51, 5, 44, 46]. Some of these works [50, 38, 47] assume knowledge of the labels of the sensitive attribute and then debias the models such that there is no correlation between the decision attribute and the sensitive attribute. IMLE-GAN [51] ensures coverage of minority groups by combining GAN adversarial training with Implicit Maximum Likelihood Estimation (IMLE) [20]. Another body of works employs a balanced unlabelled reference dataset to ensure unbiased generations [5, 44, 46]. Choi et al. [5] use a density-ratio based technique to identify the bias in datasets via the reference dataset, and learn a fair model based on importance reweighting with the help of both the original biased and the reference dataset. To capture the distance between the small reference data and the generated data, Um et al. [46] use the LeCam Divergence [19]. On the other hand, Teo et al. [44] introduce a transfer learning approach to solve this problem by training the model on the biased dataset first and then adapting the model to the reference set.
Debiasing generative models without training. As training of GANs and DMs can be resource-consuming, many methods prefer fair generation of images without explicit training [31, 43, 24, 12]. MaGNET [12] aims to produce uniform sampling on the learned manifold of any generative model like GANs or VAEs, while Ramaswamy et al.  [31] and Tan et al.  [43] manipulate the latent space of GANs to generate balanced outputs. GANDiffFace [24], on the other hand, generates balanced synthetic for face recognition by first generating high-quality images from different demographics using GANs and then finetuning Stable Diffusion [33] using DreamBooth [35] to generate more images of such identities with different poses, expressions, etc. Multiple works attempt to mitigate biases in vision-language models and text-conditioned diffusion models as well [6, 52, 41, 53, 3]. Chuang et al.  [6] debias the text embedding using a calibrated projection matrix and shows unbiased generations without any additional training or data. However, debiasing the unconditional DMs has received less attention, which is the main focus of this work.
Guidance in Diffusion Models. One of the primary techniques to condition the diffusion model is to guide the reverse diffusion model with the gradients of additional network  [8]. GLIDE [28] used CLIP [30] based guidance for open world caption to image generation. Guidance is used for image-to-image translation [49], counterfactual generation [37]. However, guidance in its original form requires retraining of guiding networks on the noisy data from scratch. Few works overcome this by learning a mapping function from the diffusion feature space to sketches for sketch-guidance  [48] and a universal guidance that repurposes pretrained networks for guidance  [1].
3 Method
We assume a setting where we are given a pretrained DM trained on biased data and a reference distribution of the sensitive attributes $\mathbf{p^{a}_{ref}}$. Our goal is to generate data from the DM, whose generated attribute distribution $\mathbf{p^{a}_{\theta}}$ best approximates the reference $\mathbf{p^{a}_{ref}}$ without retraining.
The key idea is to jointly guide the denoising of a batch of samples such that $\mathbf{p^{a}_{ref}}$approx $\mathbf{p^{a}_{\theta}}$. Directly computing $\mathbf{p^{a}_{\theta}}$ in a closed form is intractable. Instead, we train an Attribute Distribution Predictor, a small network that maps the intermediate batch features of a diffusion network to an estimate of attribute distribution $\mathbf{\hat{p}^{a}_{\theta}}$.
Figure 2: Sample guidance vs Distribution guidance. a) After a few steps of denoising ($t=\tau$), the generated samples have learned some discriminative features for gender. Sample guidance enforces a preset attribute state (male/female) to all the samples in the batch. This is stringent and can result in enforcing samples with dominant female features to be changed. However, in distribution guidance, only those samples are converted, which are easy to convert and are close to the decision boundary. b) Visualization of the generated samples with different guidance strengths. It can be observed that sample guidance is not able to change the attribute in some cases, and increasing guidance strength results in collapse.
3.1 Preliminary
Diffusion models have emerged as a powerful family of generative models trained to learn the data distribution by gradual denoising from a Gaussian distribution. Starting from a clean point $\mathbf{x_{0}}$, and a set of scalar values $\{\alpha_{t}\}_{t=1}^{T}$, applying $\mathbf{t}$ steps of the forward diffusion process yields a noisy data point, where $\bar{\alpha_{t}}$ = $\prod_{i=1}^{t}\alpha_{i}$
$\mathbf{x_{t}}=\sqrt{\bar{\alpha_{t}}}\mathbf{x_{0}}+(\sqrt{1-\bar{\alpha_{t}}})\epsilon,\epsilon\approx\mathcal{N}(0,I)$ (1)
A diffusion model is learned as a neural network $\epsilon_{\theta}$ that predicts the noise from given $\mathbf{x_{t}}$ and $\mathbf{t}$. The reverse process takes the form $q(\mathbf{x_{t-1}}|\mathbf{x_{t}},\mathbf{x_{0}})$, which is parameterized as a Gaussian distribution. In this work, we consider DDIM [42] sampling which first computes a clean data point $\mathbf{\hat{x}_{0}}$ and then sample $\mathbf{x_{t-1}}$ from $q(\mathbf{x_{t-1}}|\mathbf{x_{t}},\mathbf{\hat{x}_{0}})$.
Classifier guidance is proposed to condition a diffusion model on class labels with the help of a pretrained classifier [8]. Specifically, a classifier $f_{\phi}(c|\mathbf{x_{t}},t)$ is trained on noisy images at different noise scales, where $c$ is the class label. The gradients of the classifier are then used to guide the diffusion sampling process towards the class label $c$. Concretely, the classifier guidance performs sampling by updating the noise prediction $\epsilon_{\theta}(\mathbf{x_{t}},\mathbf{t})$.
$\hat{\epsilon}_{\theta}(\mathbf{x_{t}},t)=\epsilon_{\theta}(\mathbf{x_{t}},t)-\sqrt{1-\alpha_{t}}\nabla_{\mathbf{x_{t}}}\log f_{\phi}(c|\mathbf{x_{t}},t)$ (2)
3.2 Classifier guidance for debiasing
A promising approach is to leverage pretrained attribute classifiers to guide toward balanced generation. Assuming a reference attribute distribution $\mathbf{p^{a}_{ref}}$ for a binary attribute $\mathbf{a}$ (e.g., gender) and corresponding attribute classifier ${f}_{\phi}$. We parameterize $\mathbf{p^{a}_{ref}}$ as a Bernoulli distribution with parameter $\mathbf{r}$ denoting fraction of males samples. To generate samples following $\mathbf{p^{a}_{ref}}$, we can randomly select $\mathbf{Nr}$ samples from the batch size of $\mathbf{N}$ and guide them towards male class using predictions from ${f}_{\phi}$. Similarly, for the remaining $\mathbf{N(1-r)}$ samples, we guide them towards female class. In practice, such a sample guidance follows $\mathbf{p^{a}_{ref}}$ up to some extent, but results in inferior sample quality (see Fig. 2b).
Insight. 1: Attempting to steer the generation of individual samples towards a pre-defined attribute class is overly restrictive and leads to inferior generation quality.
Remark: Sample guidance requires enforcing a preset attribute state (male / female) to each sample of a batch during the denoising process, which is too stringent and results in distorting the outputs. Given an intermediate time-stamp $\tau$, the samples in the earlier stages ($t>\tau$) of reverse diffusion process are close to the noise space, resulting in the poor classifier of the corresponding stages. The guidance becomes effective only in the later stage of denoising ($t<\tau$). However, till timestep $\tau$, some facial features are already formed for the samples [25]. Hence, enforcing a preset attribute state at later stages ($t<\tau$) is restrictive and results in sample collapse (Fig. 2-b)). For e.g., for gender attribute, if a sample has formed dominant female features till $\tau$, enforcing it to be $male$ is overly restrictive, and classifier guidance distorts the generated sample.
3.3 Distribution Guidance
We propose an alternate guidance strategy termed Distribution Guidance for fair generation, which provides more flexibility to modify the attribute states during the denoising process. The key idea is to jointly denoise a batch of samples $\mathbf{x_{T}^{[1:N]}}$, with the guidance from reference attribute distribution $\mathbf{p^{a}_{ref}}$. To realize distribution guidance, we define a differentiable distribution prediction function $g_{\psi}$ that maps the batch samples to an estimate of generated attribute distribution $\mathbf{\hat{p}^{a}_{\theta}}$. We learn the function $g_{\psi}$ over an intermediate feature space - h-space [18] of denoising network instead of image space for efficiency (Sec.3.4). Hence, the batch estimate is given by, $\mathbf{\hat{p}^{a}_{\theta}}=g_{\psi}(\mathbf{h_{t}^{[1:N]}},\mathbf{t})$, where $\mathbf{h_{t}^{[1:N]}}$ is the bottleneck U-Net feature representation of the batch samples $\mathbf{x_{t}^{[1:N]}}$. Further, we define a loss function $\mathcal{L}(.)$ that measures the similarity of two distributions. During denoising we guide the batch of samples to bring $\mathbf{\hat{p}^{a}_{\theta}}$ closer to the reference, i.e., $\mathcal{L}(\mathbf{\hat{p}^{a}_{\theta}},\mathbf{p^{A}_{ref}})\approx 0$. This can be easily integrated into the reverse diffusion process as an extension of classifier guidance by modifying Eq.2.
Insight.2: Distribution guidance provides flexibility to batch samples to easily modify the attributes to match the defined attribute distribution.
Remark: As distribution guidance does not require presetting for attribute states for each sample, it gives more flexibility during the generation as long as $\mathbf{\hat{p}^{a}_{\theta}}$ follows $\mathbf{p^{a}_{ref}}$. E.g., in our running example of gender attribute, only those samples will change gender after $t=\tau$, which are close to the decision boundary. In contrast to sample guidance, where random samples are forced to change the attribute state (Fig.  2) resulting in inferior quality, distribution guidance is able to generate fair samples with good quality. Observe that distribution guidance covers the samples that are close to the decision boundary for gender classification, resulting in good outputs.
A major design decision for implementing distribution guidance is the selection of function $g_{\psi}$. The conventional approach is to train multiple attribute classifiers at each level of noise in the image space. Instead, we propose to leverage intermediate semantic features from the diffusion model itself and repurpose them for the task of attribute classification. This framework is super efficient as compared to image space classifiers.
3.4 Guidance in the H-space
Diffusion models, although trained for image generation, learn semantically rich representations in the internal features of denoising network $\mathbf{{\epsilon_{\theta}}}$. These representations have been successfully used for segmentation [2] and classification [26]. Motivated by this, we ask can we use these internal features for training guidance classifiers? To answer this, we take features from the bottleneck layer of the diffusion U-Net, termed as the h-space [18] and attach a linear head for classification. The trained classifiers achieve good classification performance across multiple attributes, as shown in Fig. 3. Hence, we use the trained $h-space$ classifiers to realize $g_{\psi}$. Specifically, we train a network Attribute Distribution Predictor that maps the batch h-space features to an estimate of attribute distribution $\mathbf{\hat{p}^{a}_{\theta}}$.
Figure 3: Classification accuracy for linear h-space classifiers and ResNet-18 image space classifiers trained on $2K$ training examples. h-space classifiers are data efficient and achieve superior performance even with a linear layer
Figure 4: Distribution guidance in the h-space. For a given batch $\mathbf{x_{t}^{[1:N]}}$, we extract the intermediate h-space representation $\mathbf{h_{t}^{[1:N]}}$ and pass it through ADP to obtain attribute distribution $\mathbf{\hat{p^{a}_{\theta}}}$. Guidance updates $\mathbf{h_{t}^{[1:N]}}$ by backpropagating the derivative of loss.
Attribute Distribution Predictor (ADP) is realized via a linear attribute classifier conditioned on the diffusion time step $\mathbf{t}$. Given a batch of generating samples $\mathbf{x_{t}^{[1:N]}}$, we extract the corresponding h-space features $\mathbf{h_{t}^{[1:N]}}$ = $\epsilon^{E}_{\theta}(\mathbf{x_{t}^{[1:N]}},\mathbf{t})$, where $\epsilon^{E}_{\theta}$ is the encoder of U-Net network. Next, we pass the batch $\mathbf{h_{t}^{[1:N]}}$ to the attribute classifier and obtain a batch of softmax predictions. The softmax predictions are aggregated per class to obtain the attribute distribution estimate $\mathbf{\hat{p}^{a}_{\theta}}$. Finally, we update the intermediate h-vectors with the gradients of the distribution matching loss $\mathcal{L}$, where $\mathbf{\gamma}$ is the guidance strength parameter:
$\mathbf{\tilde{h}_{t}^{[1:N]}}=\mathbf{h_{t}^{[1:N]}}$ $-\gamma*\nabla_{\mathbf{h_{t}^{[1:N]}}}\mathcal{L}(\mathbf{\hat{p}^{a}_{\theta}},\mathbf{p^{a}_{ref}})$
$\mathbf{\hat{p}^{a}_{\theta}}=$ $g_{\psi}(\mathbf{h_{t}^{[1:N]}},\mathbf{t})$
Finally, we obtain the distribution guided noise predictions $\tilde{\epsilon}$ by passing the $\mathbf{h^{[1:N]}_{t}}$ through the U-Net decoder $\epsilon^{D}_{\theta}$, i.e., $\tilde{\epsilon}(\mathbf{x^{[1:N]}_{t}},\mathbf{t})=\epsilon^{D}_{\theta}(\mathbf{h^{[1:N]}_{t}},\mathbf{t})$. The predicted noise is then used to update batch $x^{[1:N]}_{t-1}$ using DDIM [42].
Insight.3: H-space guidance is extremely effective and efficient as compared to image space guidance.
Remark. As compared to conventional image space guidance, guidance in h-space has multiple advantages: i) it requires only a set of linear layers to implement the classifiers, ii) it is fast to backpropagate during guidance as compared to image models, iii) highly data-efficient and can be trained with only a few thousand examples due to semantically rich h-space. In the experiments section, we compare these properties of the h-space guidance.
4 Experiments
In this section, we first discuss the evaluation metrics, followed by the implementation details. We take face datasets as the subject of study, as they are subject to very high demographic biases. We compare our method with other inference time debiasing approaches for both single-attribute and multi-attribute debiasing. Further, we present detailed ablations to validate design choices. Finally, we show the generalization of our debiasing approach on large text-to-image generation model Stable Diffusion [33].
4.1 Evaluation Metrics
A fair generative model is evaluated on two grounds: image quality and fairness. Therefore, we begin by discussing the metrics we use to measure these two attributes of the generated images.
Fairness. The primary goal of this paper is to generate images based on the reference distribution $\mathbf{p_{ref}^{a}}$. We follow Choi et al.  [5] to define the Fairness Discrepancy (FD) metric. Given an attribute $\mathbf{a}$, we assume access to a high-accuracy classifier for $\mathbf{a}$ (denoted as $\mathcal{C}_{a}$), and using the predictions from the latter, we compute the following [5, 44]:
$|\bar{p}-\mathbb{E}_{\mathbf{x}\sim p_{\theta}(\mathbf{x})}(\mathbf{y})|_{2}$
where $\mathbf{y}$ is the softmax output of the classifier $\mathcal{C}_{a}(\mathbf{x})$, $\bar{p}$ is a uniform vector of the same dimension as $\mathbf{y}$, $p_{\theta}$ is the distribution of the generated images. The lower the FD score, the closer the distribution of the attribute values is to the uniform distribution – i.e., the generated images are fairer with respect to attribute $\mathbf{a}$.
Image Quality. To measure the quality of generation, we follow the standard Fréchet Inception Distance (FID) [10]. We compute this metric with the help of an attribute-wise balanced dataset sampled from the original training data.
4.2 Implementation Details
Training h-space classifiers. We start with creating a paired dataset $\mathcal{D}^{h}_{clf}$ of h-vector and attribute labels. Specifically, we take a subset $\mathcal{D}$ of the CelebA-HQ [15] dataset and obtain attribute labels for the images using the pretrained attribute classifier $\mathcal{C}_{a}$. Next, we embed image $\mathbf{I^{i}}\in\mathbf{D}$ to obtain the corresponding h-space representation $\mathbf{\mathcal{H}^{i}}=\{\mathbf{h^{i}_{t}}\}^{t=T}_{t=0}$ using DDIM [42] inversion. This yields a labelled dataset $\mathcal{D}^{h}_{clf}$ with pairs $(\mathcal{H}^{i},\mathbf{y^{i}})$, where $\mathbf{y^{i}}=\mathcal{C}_{a}(I^{i})$ is the predicted attribute label for image $\mathbf{I^{i}}$ (e.g. male / female). Next, we train h-space attribute classifiers $\mathcal{C}^{h}_{a}(\mathbf{h_{t}},t)$ as a linear head over $\mathbf{h_{t}}$ and conditioned on time $t$ (ignored sample index $i$ for brevity). These obtained classifiers generate high-accuracy attribute predictions, as shown in Fig. 3. Further details about the dataset and classifiers are provided in a supplementary document.
Figure 5: Balancing generated data across gender and eyeglass attributes with various guidance strategies. Minority group $male$ and $eyeglasses$ are marked. Image-space Universal Guidance and Sample Guidance results generate imbalanced and poor-quality images. For gender, Sample Guidance is able to generate only a few males while maintaining quality. The proposed Distribution Guidance balances (close to the ratio of $0.50:0.50$) both the attributes and generates high-quality images.
Distribution guidance. To perform distribution guidance, we realize the distribution prediction function, ADP with h-space classifiers $\mathcal{C}^{h}_{a}$. Specifically for a generating batch $\mathbf{x_{t}^{[1:N]}}$, we obtain the corresponding $h-space$ representation $\mathbf{h_{t}^{[1:N]}}$ and obtain a set of attribute predictions $\mathbf{\hat{y}^{[1:N]}_{t}}$ from classifiers $\mathcal{C}^{h}_{a}$. Finally, we add all the softmax values for all the $N$ images in the batch for each class to obtain the estimate of $\mathbf{p^{a}_{\theta}}$. We use Chi-square distance as the loss $\mathcal{L}$ with the reference distribution $\mathbf{p^{a}_{ref}}$.
Diffusion Model Architecture. We evaluated our proposed debiasing method for fair generation from two state-of-the-art pretrained DMs: a) an unconditional DM, P2  [4], trained on CelebA-HQ dataset, and b) a text conditional DM StableDiffusion v1.5  [33] trained on the LAION [40] dataset. Both these models have exceptional image generation quality; however, they have significant bias concerning the sensitive face attributes, as shown in the following subsections.
4.3 Evaluation and Comparisons.
We compare our proposed method against two techniques for guidance-based generations for DMs and a state-of-the-art sampling-based technique for debiasing generative models, MagNet [12].
Sample guidance. We use classifier guidance as explained in Sec. 3.2 in the h-space using the trained h-space attribute classifiers $\mathcal{C}_{a}^{h}$. Such guidance requires presetting an attribute state for each generating image and pushing the trajectory to change the attribute state.
Universal guidance [1] performs guidance in the image space but uses pretrained classifiers trained on the clean image. This resolves the additional requirement of training image classifiers on noisy images. The key idea is to use the DDIM scheduler and predict the approximation of $\mathbf{\hat{x_{0}}}$ from noise image $\mathbf{x_{t}}$, and pass it through pretrained image classifier $\mathcal{C}_{A}$. However, this process has two shortcomings. Firstly, it is slow as we have to backpropagate the gradients from the image classifier. Experimentally, we found the h-space guidance to be $7$ times faster than the image guidance. Secondly, it performs poorly at the early stage due to an inaccurate approximation of $\mathbf{x_{0}}$. We used the same training set of $|\mathcal{D}_{clf}|=2K$ to train the h-space and image space attribute classifiers for a fair comparison.
Table 1: Evaluation of balanced generation for single attribute
Gender Race Eyeglasses Metric FD FID FD FID FD FID Random Sampling 0.178 54.59 0.334 60.01 0.251 75.21 Universal Guidance  [1] 0.193 52.10 0.377 93.42 0.189 64.55 H-Sample Guidance 0.113 51.46 0.184 56.53 0.118 57.63 H-Distribution Guidance 0.049 50.27 0.113 52.38 0.014 51.78 Random sampling* 0.307 112.28 0.463 123.97 0.276 117.83 Magnet* 0.267 91.15 0.454 97.05 0.281 106.55
Quantitative evaluation. We evaluate our debiasing method for the single attribute case by generating balanced distributions of individual sensitive attributes - gender, eyeglasses, and race, in Tab. 1. For the scope of this paper, we consider all these attributes as binary, and hence expect a $1:1$ ratio of the sensitive attributes in the synthesized images (for e.g. $0.50$ fraction of males and $0.50$ fraction of females in case of gender). Specifically, we generate $10K$ images from each method per attribute and compute the metrics defined in Sec. 4.1. The proposed guidance method outperforms all the methods in terms of visual quality as measured by FID and bias generation measured by FD. Sample guidance achieves comparable FID for gender; however, higher FD indicates inferior debiasing. The tradeoff between FD and FID is discussed in detail in subsection 4.5. This supports our thesis that the distribution guidance provides enough flexibility during generation, resulting in high-quality outputs even with a high guidance scale.
Multiple attributes. Additionally, we apply our method for debiasing multiple attributes simultaneously in Tab. 2. Specifically, given two reference distributions $\mathbf{p^{a_{1}}_{ref}}$ and $\mathbf{p^{a_{2}}_{ref}}$, we add guidance from two pretrained attribute distribution predictors $g_{\psi_{1}}$ and $g_{\psi_{2}}$. For this experiment, we define both the reference distribution as uniform for each attribute ($50\%$-$50\%$ splits). The generated results follow the reference distribution, resulting in a balanced generation across attributes.
Table 2: Evaluation of balanced generation for multiple attribute
Gender + Race Eyeglasses + Race Gender + Eyeglasses Metric FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ Random Sampling 0.256 60.68 0.292 89.14 0.214 70.97 Universal Guidance 0.283 71.84 0.264 91.54 0.157 80.57 H-Sample Guidance 0.241 59.78 0.135 67.87 0.079 52.03 H-Distribution Guidance 0.075 49.91 0.101 57.46 0.057 47.45
Comparison with MagNet [12]: MagNet is an unsupervised method enabling fair sampling from a pretrained model. They propose a method for uniform sampling on the image manifold to generate under-represented groups equally. We generated results by MagNet sampling from a StyleGAN2 [17] model trained on the FFHQ [16] dataset from their official codebase. We compare our method with MagNet [12] for the task of fair generation in Tab. 1. We used MagNet to generate $10K$ samples and compute the FD score. As the base model for MagNet is StyleGAN2, we cannot compare FID directly with our DM debiased results. However, we can observe that MagNet achieves similar FD scores to its baseline model and is ineffective in debiasing the three sensitive attributes: gender, race, and eyeglasses. However, our distribution guidance significantly improves the FD metric from the baseline DM model.
Qualitative evaluation. We present qualitative results for attribute balancing with respect to gender and eyeglasses in Fig. 5. We randomly sample $20$ starting noise and use individual guidance methods for debiasing. Without guidance, the DM mostly generates female faces for the gender attribute. Sample guidance and universal guidance, although increasing the number of males, collapses some images. On the other hand, Distribution guidance generates an almost equal number of males and females without affecting the generation quality. Moreover, the DM hardly generates images with eyeglasses for sample and universal guidance and also without any guidance. Our method, on the contrary, leads to highly balanced generations.
4.4 Generating Imbalanced Distributions
We tested our distribution guidance for generating imbalanced attribute distribution by providing skewed $\mathbf{p^{a}_{ref}}$ i) $0.20$ female and $0.80$ male ii) $0.10$ white race and $0.90$ black race. These two settings are extremely challenging given that male and black race are minority groups in the dataset. We present qualitative results for both the settings in Fig.6, where our distribution guidance is able to generate the defined distribution with majority males. Note, we have binarized the race attribute as black and whites for simplicity, and hence in the generation brown race is considered under black category. We report quantitative metrics in Tab. 3, where the proposed method can achieve good FD scores with the same FID.
Figure 6: Generating imbalanced data with two non-uniform reference distribution - i) 0.9B-0.1W, ii) 0.8M-0.2F. Minority classes are marked. Distribution guidance generates high-quality images and closely follows the skewed reference distribution.
Table 3: Distribution guidance for imbalanced generation
0.20F - 0.80M 0.10W - 0.90B Method FD FID FD FID Random Sampling 0.478 72.26 0.734 77.63 H-Distribution Sampling 0.168 51.65 0.325 53.80
4.5 Ablations
Guidance Strength $\gamma$. We ablate over guidance strength parameter $\gamma$ for both the distribution guidance and sample guidance in Fig. 7. Increasing $\gamma$ reduces the bias (better FD) at the cost of inferior image quality (increased FID). Guidance strength $1500$ achieved a good tradeoff between image quality and balancing attributes. Further, as compared to sample guidance, the distribution guidance achieves a better tradeoff for all the guidance strengths.
Figure 7: Ablation over guidance strength $\gamma$
Batch size. As we approximate $\mathbf{p^{a}_{\theta}}$ with a an estimate $\mathbf{\hat{p}^{a}_{\theta}}$ over a batch of size $N$, we ablate over different values of $N$ in Tab. 4. Intuitively, using a larger batch size yields a better estimate of $\mathbf{p^{a}_{\theta}}$. We have found that the $N=100$ works best for all our experiments.
Table 4: Ablation over batch size for gender balancing.
Batch size 10 25 50 75 100 FD $\downarrow$ 0.062 0.059 0.049 0.059 0.046 FID $\downarrow$ 50.54 51.51 50.86 50.98 51.64
Figure 8: Debiasing results on stable diffusion for gender attribute. The proposed distribution guidance can balance the gender attributes in the neutral prompt and generalizes well to prompts involving other professions e.g. firefighter, doctor.
4.6 Experiments on Stable Diffusion
We implemented our distribution guidance technique for debiasing a text-to-image generation model, Stable Diffusion (SD) v1.5 [33] concerning the gender attribute in this subsection. First, we generate a dataset $\mathcal{D}$, of $10K$ images from SD, with prompts ‘a photo of a male’ and ‘a photo of a female’ to generate a labeled dataset for training h-space classifiers. If we use the same CelebA-HQ dataset to train the h-space classifier, the guidance is ineffective due to a significant domain shift from the SD generations. Next, we obtain the corresponding labeled dataset $\mathcal{D}^{h}_{clf}$ in the h-space and train a gender classifier in the h-space. We used the trained classifier for distribution guidance following Sec. 3.4 to generate images with balanced gender.
It is observed that SD increases the gender bias when queried to generate certain professions (e.g. male and doctor) [52]. To this end, we implement our distribution guidance along the prompts ‘a photo of a doctor’ and ‘a photo of a firefighter’ to evaluate the effectiveness in this challenging setting. The qualitative results are shown in Fig.8, and the quantitative results are reported in Tab. 5. We can observe, as compared to the baseline (random generation without guidance), our method generates fairer gender distribution with both neutral prompts and prompts mentioning professions (e.g. ‘a photo of a firefighter’). Additional results on text-to-image models are present in supp. material.
Table 5: Balancing gender on Stable Diffusion [33] model.
Method Gender Doctor Firefighter FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ Random Sampling 0.317 72.37 0.355 70.11 0.235 71.86 H-Sample Guidance 0.026 70.96 0.021 68.43 0.097 70.42 H-Distribution Guidance 0.024 70.69 0.015 67.36 0.093 69.41
4.7 Class-imbalance in attribute classification
We explore an important downstream application of our proposed approach in balancing minority classes by augmenting the under-represented classes with generated data. Specifically, we train a race classifier (labels obtained from $\mathcal{C}_{\mathcal{A}}$) on the CelebA-HQ [14] dataset. The race classifier is an ImageNet [7]-pretrained ResNet-18 [9] encoder, followed by 2 MLP layers and a classifier. We manually oversample the Whites and undersample the Blacks in the training dataset such that the imbalanced dataset consists of $10k$ samples of white people and $1k$ black people (we keep the genders balanced within a race class). Consequently, the model performs poorly on the minority class (i.e., Black) due to under-representation. Next, we augment (class-balanced) the training data by generating samples whose distribution is inversely proportional to the class counts in the training set to increase images of minority classes using our distribution guidance approach. This adds 9k images of only Black race, and the classifier trained on this balanced data performs significantly better (Tab. 6). Even when gender is balanced in both the classes, we observe a significant disparity in the accuracies for Male and Female Black samples in the vanilla classifier. However, our proposed method helps reduce the accuracy gap between Black males and Black females. This shows a potential application in generating ’class’ balanced datasets to train models for other downstream tasks, which can also mitigate bias.
Table 6: Group-wise accuracies of identical classifiers trained on existing and generated(balanced) data
Black female White female Black male White male Vanilla classifier 75.76 98.96 70.29 99.93 Balanced classifier 91.04 97.33 90.72 97.68
5 Discussion.
Limitations. Although our method performs guidance in the h-space, which is efficient compared to the image space guidance, it still requires additional training of h-space classifiers. Another limitation is reliance on accurate attribute classifiers to obtain labels for training h-space classifiers.
Future works. An important future work is extending distribution guidance beyond de-biasing for controlled generation and data augmentation. In the context of debiasing DMs, extending the proposed approach without needing an attribute classifier or labeled data.
6 Conclusion
In this work, we aim to mitigate biases from pretrained diffusion models without retraining - given only a desired reference attribute distribution. We propose a novel approach leveraging distribution guidance that jointly guides a batch of images to follow the reference attribute distribution. The proposed method is effective and results in both high-quality and fair generations across multiple attributes and outperforms sample guidance strategies based on conditioning each sample individually. Extensive experiments demonstrate the effectiveness of our method in balancing both single and multiple attributes on unconditional DMs and conditional text-to-image diffusion models. We believe such a setting of debiasing without retraining is practical, especially in today’s era of large-scale generative models.